# Generative AI for Cybersecurity of Energy Management Systems: Methods, Challenges, and Future Directions

### v1

@@ 3.3.1 A Normal HMI Screen with Power Flow Analysis Using Only GenAI Without SoM Technique
Proposed Set-of-Mark Guided Generative Intelligence (SoM-GI) Technique
The arrangement of power system segments into structured grid layouts presents significant challenges in visualization and analysis in terms of different unexpected scenarios that can happen in the SCADA room. In this section, adapting the SoM prompting methodology proposed, combined with the GenAI (named as SoM-GI) to systematically arrange different segments based on visual markers and linguistic rules, including CBs, directional indicators, and CPs, and semantic understanding of the GPT model.
Problem Formulation
Following the SoM-GI framework, the power system arrangement task is defined as a visual grounding problem. Given a set of power system segments $\mathcal{S}=\{S_{1},S_{2},...,S_{9}\}$ to be arranged in a $3\times 3$ grid, each segment contains visual markers that guide proper placement. Let $G\in\mathbb{R}^{3\times 3}$ represent the target grid arrangement, where each cell $G_{i,j}$ contains one segment from $\mathcal{S}$. The arrangement function is defined as Eq. (21[46]:
@@
where $\mathcal{M}$ represents the set of visual markers including CBs, directional indicators, and CPs.
Power Flow Principles
PF analysis constitutes a fundamental component of SCADA systems, encompassing tasks of monitoring, planning, and operational control. In the context of PF analysis, an SCADA system is typically employed to determine the electrical network’s structure, where buses serve as pivotal nodes within the diagram. An SLD illustrates the interconnections between different buses through transmission lines that quantify the power being transferred from one bus to another. Additionally, the SLD may highlight which generators are integrated into the automatic generation control (AGC) system, a mechanism that balances power supply and demand by regulating the output of generators [48, 49]. Different CBs, current flows, impedances, and transmission lines are other components of SLDs. This visual evaluation is necessary and crucial in SCADA rooms of utilities to monitor the operation of the system for potential errors and, thus, take appropriate actions to restore them when faced with outages/interruptions [50, 51]. These SLDs consist of multiple buses, interconnected by CBs (e.g., $CB_{i}$ and $CB_{ij}$) such that each bus may have associated loads or distributed generation units. Also, they include other components such as generators, transmission lines, impedances, and current flows. As an example, $CB_{ij}$ illustrates the CB between buses $i$ and $j$, near bus $i$, and all CBs operate in their normal conditions (black color), as demonstrated in Fig. 7.
@@
The abnormal scenario is based on the $CB_{ji}$ failure due to communication errors which prevent the fault isolation in the transmission line between buses $i$ and $j$. According to this condition, the rest of the network can work with other DERs. Additionally, these anomalous behaviors can be observed: $CB_{ji}$ remains closed visualized with a black square. A communication/SCADA error prevents the open command from reaching $CB_{ji}$; hence, the fault persists, causing a power outage in the downstream network.
The formulations and rules to train the GPT model are defined in a comprehensive way suitable for the PF analysis that a part of these principles are given as Eqs. (22)–(41[52, 53]. These principles are specifically defined in this application to analyze the SCADA information by training the GPT model.
@@
$P_{Bij}^{(t)}=\frac{|V_{Bi}|^{2}R_{ij}-|V_{Bi}||V_{Bj}|\cos(\theta_{i}-\theta_{j})R_{ij}+|V_{Bi}||V_{Bj}|\sin(\theta_{i}-\theta_{j})X_{ij}}{R_{ij}^{2}+X_{ij}^{2}}$ (30)
$Q_{Bij}^{(t)}=\frac{|V_{Bi}|^{2}X_{ij}-|V_{Bi}||V_{Bj}|\cos(\theta_{i}-\theta_{j})X_{ij}-|V_{Bi}||V_{Bj}|\sin(\theta_{i}-\theta_{j})R_{ij}}{R_{ij}^{2}+X_{ij}^{2}}$ (31)
$P_{Bji}^{(t)}=\frac{|V_{Bj}|^{2}R_{ij}-|V_{Bj}||V_{Bi}|\cos(\theta_{j}-\theta_{i})R_{ij}+|V_{Bj}||V_{Bi}|\sin(\theta_{j}-\theta_{i})X_{ij}}{R_{ij}^{2}+X_{ij}^{2}}$ (32)
$Q_{Bji}^{(t)}=\frac{|V_{Bj}|^{2}X_{ij}-|V_{Bj}||V_{Bi}|\cos(\theta_{j}-\theta_{i})X_{ij}-|V_{Bj}||V_{Bi}|\sin(\theta_{j}-\theta_{i})R_{ij}}{R_{ij}^{2}+X_{ij}^{2}}$ (33)
$P_{Bjk}^{(t)}=\frac{|V_{Bj}|^{2}R_{jk}-|V_{Bj}||V_{Bk}|\cos(\theta_{j}-\theta_{k})R_{jk}+|V_{Bj}||V_{Bk}|\sin(\theta_{j}-\theta_{k})X_{jk}}{R_{jk}^{2}+X_{jk}^{2}}$ (34)
$Q_{Bjk}^{(t)}=\frac{|V_{Bj}|^{2}X_{jk}-|V_{Bj}||V_{Bk}|\cos(\theta_{j}-\theta_{k})X_{jk}-|V_{Bj}||V_{Bk}|\sin(\theta_{j}-\theta_{k})R_{jk}}{R_{jk}^{2}+X_{jk}^{2}}$ (35)
$P_{Bkj}^{(t)}=\frac{|V_{Bk}|^{2}R_{jk}-|V_{Bk}||V_{Bj}|\cos(\theta_{k}-\theta_{j})R_{jk}+|V_{Bk}||V_{Bj}|\sin(\theta_{k}-\theta_{j})X_{jk}}{R_{jk}^{2}+X_{jk}^{2}}$ (36)
$Q_{Bkj}^{(t)}=\frac{|V_{Bk}|^{2}X_{jk}-|V_{Bk}||V_{Bj}|\cos(\theta_{k}-\theta_{j})X_{jk}-|V_{Bk}||V_{Bj}|\sin(\theta_{k}-\theta_{j})R_{jk}}{R_{jk}^{2}+X_{jk}^{2}}$ (37)
The statuses of different CBs are presented in Eqs. (38)–(41), in which the principles of CBs are assessed in opened and/or closed conditions. $CB_{ij}^{(t)}$ depicts the CB between buses $i$ and $j$, near to bus $i$. If at least one of the CBs (e.g., $CB_{ij}^{(t)}$ or $CB_{ji}^{(t)}$) is opened according to Eq. (38), there are no electric current and PFs between buses, which shows a normal condition.
@@
This similar analysis can be expanded to Eq. (39) that at least one of the CBs between buses $j$ and $k$ is opened and this part experiences an open circuit status.
@@
Please note that there are more principles according to the PF analysis which are formulated and trained according to the proposed framework during the implementation process, considering the different normal and abnormal scenarios which can happen for components such as other CB statuses during a fault in different transmission lines, voltage violations, and the presence of loads in different buses.
Circuit Breaker Markers:
CBs serve as primary connection indicators between buses at the terminals of transmission lines. For buses $i$ and $j$, the CB markers follow the notation $\text{CB}_{ij}\leftrightarrow\text{CB}_{ji}$, indicating terminal connections of transmission lines. These CBs (i.e., red square markers) establish bidirectional connectivity constraints.
Directional Indicators:
Each transmission line includes directional markers $\text{L}i\_j\_\text{d}$ where $d\in\{N,S,E,W,NE,NW,SE,SW\}$ represents the different directions. The directional constraint is:
@@
Connection Point Matching:
CPs follow a pairing scheme as $\text{CP}_{i\_j\_A}$ $\leftrightarrow$ $\text{CP}_{i\_j\_B}$ and $\text{CP}_{i\_j\_C}$ $\leftrightarrow$ $\text{CP}_{i\_j\_D}$ that these blue boundary markers ensure proper alignment of adjacent segments.
Segment Arrangement Algorithm
The arrangement process follows a constraint satisfaction approach:
@@
Implementation Approach
The SoM-GI methodology enhances visual grounding by making implicit connections explicit through markers. For power system segments:
Marker Visibility:
Each marker type uses distinct visual encoding:
@@
Spatial Reasoning:
The arrangement leverages spatial relationships encoded in directional markers. For instance, $\text{L}1\_2\_\text{S}$ indicates bus 1 connects to bus 2 via the southern boundary, requiring the segment containing bus 2 to be placed south of the segment containing bus 1. A sample of an HMI screen segment with and without the SoM-GI method is represented in Fig. 8. As can be seen, there are different indicators to enhance the understanding of the PF analysis in an HMI screen according to the transmission lines, CPs at edges/boundaries, CB markers, and load markers. The similar SoM-GI approach can be applied to other segments.
@@
Finally, according to the provided information, the following recommendations/rules are crafted to train the GPT model in addition to the technical parts of PF analysis, as shown in the following box: Recommendations/Rules for training the GPT model Red squares show the closed circuit breakers (CBs) at the terminals of each transmission line. For example, $\text{CB}1\_2$ and $\text{CB}2\_1$ should be at the terminals of one transmission line between bus 1 and bus 2. This rule applies to all buses with relevant CBs. Also, green squares show the open CBs so there is no flow of power. Also, it is not possible to have different statuses of CBs at both terminals of one single transmission line. That means that if we have a closed CB at one terminal, there should be another closed CB at another terminal to meet the normal scenario. Otherwise, this is an anomaly. Indicators $\text{L}i\_j\_\text{direction}$” show the connection of different transmission lines in which N, S, W, E, NE, NW, SE, SW show North, South, West, East, Northeast, Northwest, Southeast, Southwest, respectively. These directions show the connection to the adjacent segment according to the transmission line connections. For instance, $\text{L}1\_2\_\text{S}$ (connection of bus 1 to bus 2 through south direction) and/or $\text{L}2\_1\_\text{N}$ show an adjacent of relevant segments in a proper way. This rule applies to all transmission lines and relevant buses. There are some CP at end of some parts for better clarification. Actually, CPs are boundary points in blue for better clarification. For instance, $\text{CP}1\_2\_\text{A}$” at one edge and $\text{CP}1\_2\_\text{B}$” at another edge. It means these two points should have connected each other. Also, in some case, there are $\text{CP}i\_j\_\text{C}$” and $\text{CP}i\_j\_\text{D}$” in which “C” and “D” should connect each other. That means that “A” should be connected to “B” and “C” should be connected to “D”. Please check this rule carefully. Match connection pairs: A connects to B C connects to D Same $\text{CP}i\_j$ must connect Arrange segments so matched CPs are adjacent. This approach provides several advantages for power system visualization, including:
@@
Based on these connections, the $3\times 3$ grid arrangement should be:
@@
3.3.2 Scenario #3B: A CB Malfunction Between Bus 6 and Bus 13
@@
3.3.3 Scenario #3C: False Data Injection at Bus 2
@@
4 Conclusions and Future Directions

### v2

@@ III-C1 A Normal HMI Screen with Power Flow Analysis Using Only GenAI Without SoM Technique
Proposed Set-of-Mark Generative Intelligence (SoM-GI) Technique
The arrangement of power system segments into structured grid layouts presents significant challenges in visualization and analysis in terms of different unexpected scenarios that can happen in the SCADA room. In this section, adapting the SoM prompting methodology proposed, combined with the GenAI (named as SoM-GI) to systematically arrange different segments based on visual markers and linguistic rules, including CBs, directional indicators, and CPs, and the semantic understanding of the GPT model.
Problem Formulation Following the SoM-GI framework, the power system arrangement task is defined as a visual grounding problem. Given a set of power system segments $\mathcal{S}=\{S_{1},S_{2},...,S_{9}\}$ to be arranged in a $3\times 3$ grid, each segment contains visual markers that guide proper placement. Let $G\in\mathbb{R}^{3\times 3}$ represent the target grid arrangement, where each cell $G_{i,j}$ contains one segment from $\mathcal{S}$. The arrangement function is defined as Eq. (21[54]:
@@
where $\mathcal{M}$ represents the set of visual markers, including CBs, directional indicators, and CPs. Power Flow Principles PF analysis constitutes a fundamental component of SCADA systems, encompassing tasks of monitoring, planning, and operational control. In the context of PF analysis, an SCADA system is typically employed to determine the electrical network’s structure, where buses serve as pivotal nodes within the diagram. An SLD illustrates the interconnections between different buses through transmission lines that quantify the power being transferred from one bus to another. Additionally, the SLD may highlight which generators are integrated into the automatic generation control (AGC) system, a mechanism that balances power supply and demand by regulating the output of generators [56, 57]. Different CBs, current flows, impedances, and transmission lines are other components of SLDs. A visual assessment is essential in SCADA rooms of utilities to monitor system operations, identify potential errors, and implement corrective measures during outages or interruptions [58, 59]. These SLDs consist of multiple buses, interconnected by CBs (e.g., $CB_{i}$ and $CB_{ij}$) such that each bus may have associated loads or distributed generation units. Also, they include other components such as generators, transmission lines, impedances, and current flows. As an example, $CB_{ij}$ illustrates the CB between buses $i$ and $j$, near bus $i$, and all CBs operate in their normal conditions (black color), as demonstrated in Fig. 7.
@@
The abnormal scenario is based on the $CB_{ji}$ failure due to communication errors, which prevent the fault isolation in the transmission line between buses $i$ and $j$. According to this condition, the rest of the network can work with other DERs. Additionally, these anomalous behaviors can be observed: $CB_{ji}$ remains closed, visualized with a black square. A communication/SCADA error prevents the open command from reaching $CB_{ji}$; hence, the fault persists, causing a power outage in the downstream network. The formulations and rules to train the GPT model are defined in a comprehensive way suitable for the PF analysis that a part of these principles are given as Eqs. (22)–(41[60, 61]. These principles are specifically defined in this application to analyze the SCADA information by training the GPT model.
@@
$P_{Bij}^{(t)}=\frac{|V_{Bi}|^{2}R_{ij}-|V_{Bi}||V_{Bj}|\cos(\theta_{i}-\theta_{j})R_{ij}+|V_{Bi}||V_{Bj}|\sin(\theta_{i}-\theta_{j})X_{ij}}{R_{ij}^{2}+X_{ij}^{2}}$
$Q_{Bij}^{(t)}=\frac{|V_{Bi}|^{2}X_{ij}-|V_{Bi}||V_{Bj}|\cos(\theta_{i}-\theta_{j})X_{ij}-|V_{Bi}||V_{Bj}|\sin(\theta_{i}-\theta_{j})R_{ij}}{R_{ij}^{2}+X_{ij}^{2}}$
$P_{Bji}^{(t)}=\frac{|V_{Bj}|^{2}R_{ij}-|V_{Bj}||V_{Bi}|\cos(\theta_{j}-\theta_{i})R_{ij}+|V_{Bj}||V_{Bi}|\sin(\theta_{j}-\theta_{i})X_{ij}}{R_{ij}^{2}+X_{ij}^{2}}$
$Q_{Bji}^{(t)}=\frac{|V_{Bj}|^{2}X_{ij}-|V_{Bj}||V_{Bi}|\cos(\theta_{j}-\theta_{i})X_{ij}-|V_{Bj}||V_{Bi}|\sin(\theta_{j}-\theta_{i})R_{ij}}{R_{ij}^{2}+X_{ij}^{2}}$
$P_{Bjk}^{(t)}=\frac{|V_{Bj}|^{2}R_{jk}-|V_{Bj}||V_{Bk}|\cos(\theta_{j}-\theta_{k})R_{jk}+|V_{Bj}||V_{Bk}|\sin(\theta_{j}-\theta_{k})X_{jk}}{R_{jk}^{2}+X_{jk}^{2}}$
$Q_{Bjk}^{(t)}=\frac{|V_{Bj}|^{2}X_{jk}-|V_{Bj}||V_{Bk}|\cos(\theta_{j}-\theta_{k})X_{jk}-|V_{Bj}||V_{Bk}|\sin(\theta_{j}-\theta_{k})R_{jk}}{R_{jk}^{2}+X_{jk}^{2}}$
$P_{Bkj}^{(t)}=\frac{|V_{Bk}|^{2}R_{jk}-|V_{Bk}||V_{Bj}|\cos(\theta_{k}-\theta_{j})R_{jk}+|V_{Bk}||V_{Bj}|\sin(\theta_{k}-\theta_{j})X_{jk}}{R_{jk}^{2}+X_{jk}^{2}}$
$Q_{Bkj}^{(t)}=\frac{|V_{Bk}|^{2}X_{jk}-|V_{Bk}||V_{Bj}|\cos(\theta_{k}-\theta_{j})X_{jk}-|V_{Bk}||V_{Bj}|\sin(\theta_{k}-\theta_{j})R_{jk}}{R_{jk}^{2}+X_{jk}^{2}}$
The various CB statuses are shown in Eqs. (38)–(41), where the principles of CBs are evaluated under both open and closed conditions. $CB_{ij}^{(t)}$ depicts the CB between buses $i$ and $j$, near to bus $i$. If at least one of the CBs (e.g., $CB_{ij}^{(t)}$ or $CB_{ji}^{(t)}$) is opened according to Eq. (38), there is no electric current and PFs between buses, which show a normal condition.
@@
This similar analysis can be expanded to Eq. (39) that at least one of the CBs between buses $j$ and $k$ is opened, and this part experiences an open circuit status.
@@
Please note that there are more principles according to the PF analysis which are formulated and trained according to the proposed framework during the implementation process, considering the different normal and abnormal scenarios which can happen for components such as other CB statuses during a fault in different transmission lines, voltage violations, and the presence of loads in different buses. Circuit Breaker Markers: CBs serve as primary connection indicators between buses at the terminals of transmission lines. For buses $i$ and $j$, the CB markers follow the notation $\text{CB}_{ij}\leftrightarrow\text{CB}_{ji}$, indicating terminal connections of transmission lines. These CBs (i.e., red square markers) establish bidirectional connectivity constraints. Directional Indicators: Each transmission line includes directional markers $\text{L}i\_j\_\text{d}$ where $d\in\{N,S,E,W,NE,NW,SE,SW\}$ represents the different directions. The directional constraint is:
@@
Connection Point Matching: CPs follow a pairing scheme as $\text{CP}_{i\_j\_A}$ $\leftrightarrow$ $\text{CP}_{i\_j\_B}$ and $\text{CP}_{i\_j\_C}$ $\leftrightarrow$ $\text{CP}_{i\_j\_D}$ that these blue boundary markers ensure proper alignment of adjacent segments.
Segment Arrangement Algorithm The arrangement process follows a constraint satisfaction approach:
@@
Implementation Approach The SoM-GI methodology enhances visual grounding by making implicit connections explicit through markers. For power system segments:
Marker Visibility: Each type of marker employs unique visual encoding as follows:
@@
Spatial Reasoning: The arrangement leverages spatial relationships encoded in directional markers. For instance, $\text{L}1\_2\_\text{S}$ indicates bus 1 connects to bus 2 via the southern boundary, requiring the segment containing bus 2 to be placed south of the segment containing bus 1. A sample of an HMI screen segment with and without the SoM-GI method is represented in Fig. 8. As can be seen, there are different indicators to enhance the understanding of the PF analysis in an HMI screen according to the transmission lines, CPs at edges/boundaries, CB markers, and load markers. The analogous SoM-GI methodology is applicable and can be extended to be utilized across a diverse range of sectors and segments, allowing for a comprehensive and interdisciplinary approach within various fields of study.
@@
Finally, according to the provided information, the following recommendations/rules are crafted to train the GPT model in addition to the technical parts of PF analysis, as shown in the following box: Recommendations/Rules for the training process Red squares show the closed circuit breakers (CBs) at the terminals of each transmission line. For example, $\text{CB}1\_2$ and $\text{CB}2\_1$ should be at the terminals of one transmission line between bus 1 and bus 2. This rule applies to all buses with relevant CBs. Also, green squares show the open CBs so there is no flow of power. Also, it is not possible to have different statuses of CBs at both terminals of one single transmission line. That means that if we have a closed CB at one terminal, there should be another closed CB at another terminal to meet the normal scenario. Otherwise, this is an anomaly. Indicators $\text{L}i\_j\_\text{direction}$” show the connection of different transmission lines in which N, S, W, E, NE, NW, SE, SW show North, South, West, East, Northeast, Northwest, Southeast, Southwest, respectively. These directions show the connection to the adjacent segment according to the transmission line connections. For instance, $\text{L}1\_2\_\text{S}$ (connection of bus 1 to bus 2 through south direction) and/or $\text{L}2\_1\_\text{N}$ show an adjacent of relevant segments in a proper way. This rule applies to all transmission lines and relevant buses. There are some CP at end of some parts for better clarification. Actually, CPs are boundary points in blue for better clarification. For instance, $\text{CP}1\_2\_\text{A}$” at one edge and $\text{CP}1\_2\_\text{B}$” at another edge. It means these two points should have connected each other. Also, in some case, there are $\text{CP}i\_j\_\text{C}$” and $\text{CP}i\_j\_\text{D}$” in which “C” and “D” should connect each other. That means that “A” should be connected to “B” and “C” should be connected to “D”. Please check this rule carefully. Match connection pairs: A connects to B C connects to D Same $\text{CP}i\_j$ must connect Arrange segments so matched CPs are adjacent. This approach provides several advantages for power system visualization, including:
@@
Based on these connections, the $\times3$ grid arrangement should be:
@@
III-C2 Scenario #3B: A CB Malfunction Between Bus 6 and Bus 13
@@
III-C3 Scenario #3C: False Data Injection at Bus 2
@@
III-D LIMITATIONS AND CONSIDERATIONS
The GenAI-based framework proposed for enhancing cybersecurity in EMSs demonstrates significant potential, yet it is crucial to recognize several key limitations for a comprehensive assessment. Initially, the existing system relies on cloud-based API interactions with Anthropic Claude Pro, causing delays in response times. These delays are inadequate for real-time control tasks requiring responses within sub-second intervals. To overcome this constraint, it is probable that edge-computing systems or locally deployed AI models will need to be utilized to satisfy demanding temporal constraints.
Secondly, the framework’s reliance on external cloud services could present vulnerabilities, such as dependence on service availability, risk of sensitive grid data exposure, ongoing subscription expenses, and restricted authority over model updates or retraining timelines. Thirdly, the system’s performance is significantly dependent on the quality, quantity, and diversity of its training data. Despite the fact that validation was conducted utilizing different scenarios on the IEEE 14-bus test system, additional comprehensive evaluation is imperative to establish both scalability and robustness of the approach, particularly when it is deployed on larger grid systems or when encountering unforeseen attack scenarios that were not part of the initial validation set. Models optimized for one setup may fail to generalize well to different configurations without re-optimization or retraining. As systems expand, scalability becomes a significant challenge, especially with the increasing incorporation of distributed energy resources (DERs). The SoM-GI methodology may experience exponential computational complexity with larger HMI displays, and the processing of high-frequency PMU data streams will require additional algorithmic optimization. In addition, while GenAI offers understandable explanations in natural language, its somewhat ambiguous reasoning can lead to trust issues for operators familiar with deterministic approaches. This matter is especially crucial in safety-critical settings, where AI tools should function as decision-support mechanisms rather than independent decision entities. A significant challenge lies in regulatory compliance. The existing NERC CIP standards provide insufficient direction for the certification, validation, and auditing processes of AI-based systems, potentially hindering their deployment in regulated utility contexts. The validation performed with simulated IEEE 14-bus data, though controlled and systematic, fails to entirely encompass the complexities found in real-world power systems. These complexities include factors such as sensor noise, communication delays, and measurement uncertainties. Consequently, there is a need for future field testing under actual operational conditions. Ultimately, the costs associated with implementation including computing infrastructure, software licenses, staff training, and maintenance must be thoroughly evaluated. These expenses could lead to discrepancies, as only large utilities might be able to afford sophisticated AI-based defense systems. Despite these limitations, the proposed GenAI-based framework marks a significant step forward in the cybersecurity of modern power systems. When used as an additional layer in a defense-in-depth strategy, it has the potential to significantly enhance traditional protection systems and human expertise, leading to more robust and adaptive grid security.
IV Conclusions and Future Directions